# Correlation Functions and Chaotic Behavior of the SYK Chain Model in Pure States

### v1

@@ Abstract
Recent investigations of Rényi entanglement entropy in the SYK chain of Majorana fermions have indicated that the model exhibits slow thermalization when initialized in certain states. The extent to which the heavy modes–believed to underlie this behavior–affect other aspects of the model remains an open question. In this work, I study thermalization and scrambling of information in individual energy eigenstates of the SYK chain using exact diagonalization. I show that two-point correlation functions in finite-energy eigenstates closely match their thermal counterparts and that information scrambling occurs efficiently within these pure states. These results suggest that the slow thermalization observed in entanglement dynamics does not extend to all probes of thermalization and scrambling, even in pure states. I discuss the implications of these findings for thermal states in a potential holographic dual theory.
@@ I Introduction
Chaotic quantum many-body systems are generally expected to thermalize rapidly, even in the absence of external reservoirs. The eigenstate thermalization hypothesis (ETH) provides a framework for understanding this phenomenon [1, 2, 3]. ETH posits that for generic observables, the expectation values in individual energy eigenstates are consistent with thermal averages, implying that a pure state far from equilibrium will evolve such that its observables approach and remain close to thermal values for most of the time.
@@
The SYK chain is a one-dimensional lattice model with quenched disorder, in which each site hosts $N$ Majorana fermions subject to random four-fermion interactions [5]. Neighboring sites are coupled via random bilinear interactions. This model inherits several important features from the original, zero-dimensional SYK model [7, 8], including local criticality, an extensive zero-temperature entropy, and maximal chaos. It also displays additional features such as diffusive energy transport and a butterfly velocity that characterizes the spatial spread of chaos [5]. From a condensed matter perspective, the SYK chain is a rare example of a solvable, strongly interacting, chaotic lattice model–offering a tractable platform for studying thermalization, entanglement propagation, many-body localization, and quantum transport.
From the viewpoint of holography, the SYK chain has been conjectured to be dual to an incoherent black hole [5]. Since the SYK chain is a closed, unitary system, understanding its thermalization dynamics may yield insights into black hole formation and evaporation, and thereby inform broader discussions surrounding the black hole information paradox [9].  [^1] For recent developments in resolving the information paradox, see [10, 11, 12, 13, 14, 15, 16, 17, 18].
@@ II Background
The SYK chain model [5] is a higher-dimensional generalization of the original Sachdev-Ye-Kitaev (SYK) model [7, 8]. In this section, I first provide a brief overview of the original SYK model before describing the structure and properties of the SYK chain model.
@@
To introduce spatial locality, Gu et al. [5] proposed a $(1+1)$-dimensional generalization, known as the SYK chain model. This model consists of a one-dimensional lattice of $M$ SYK sites (also referred to as “links”), with each site containing $N$ Majorana fermions and interacting with its nearest neighbors via quartic interactions. The structure is illustrated in Fig. 1.
@@ III Computational Setup
In my earlier work [6], I examined the eigenstate thermalization hypothesis (ETH) for the SYK chain model with $M=2$, $3$, and $4$. Here, I focus on the case $M=4$ to investigate additional indicators of thermalization and quantum information scrambling.
@@ IV Correlation Functions and Signatures of Quantum Chaos
While many of the correlation functions considered here have been previously analyzed in the thermal ensemble, my emphasis is on their behavior in pure states, particularly in individual energy eigenstates. From the perspective of the holographic correspondence, this translates into studying how accurately correlation functions computed in an incoherent black-hole background are approximated by correlations in pure states.
@@ IV.1 The Spectral Form Factor
Figure 2 displays ${\cal S}(\beta,t)$ as a function of time for the SYK chain, averaged over 1000 disorder realizations and evaluated at inverse temperature $\beta=1.5$. The data reveals the familiar three-stage structure also observed in RMT: an initial decay (the “dip”), followed by a linear ramp, and eventual saturation at a plateau [24, 22]. While this qualitative structure resembles RMT behavior, the quantitative features–such as the dip time and plateau time–depend on model-specific details [24].
@@ IV.2 Two-Point Correlation Function in Energy Eigenstates
Figure 3 compares $G^{n}_{c}(t)$ to $G^{\beta}(t)$ for various coupling regimes. Agreement is evident at both early and late times, while small discrepancies arise at intermediate times, particularly near the end of the power-law decay regime. There, $G^{n}_{c}(t)$ exhibits residual oscillations around its long-time plateau, in contrast to the smoother approach seen in $G^{\beta}(t)$. These differences are suppressed with increasing Hilbert space dimension and are expected to vanish as $N\to\infty$. This thermodynamic limit is of special relevance to studies of the putative gravitational dual, where it corresponds to the semiclassical regime.
Figure 3: Two-point correlation functions of the hopping operator between two sites of the SYK chain with $NM=24$ Majorana fermions. Each correlation function is averaged over $1000$ realizations of the disorder. The eigenstate correlators $G^{n}_{c}(t)$ and the canonical correlators $G^{\beta}(t)$ exhibit excellent agreement across all three choices of intra- and inter-site coupling strengths.
@@ IV.3 Four-Point Correlation Function in Energy Eigenstates
To match the thermal and eigenstate energies, I fix the inverse temperature to $\beta=1.5$ and select the eigenstate $|n\rangle$ whose energy satisfies
$E_{n}\sim E(\beta)=\frac{1}{Z}\text{Tr}[e^{-\beta H}H],$ (33)
where $Z=\text{Tr}[e^{-\beta H}]$ is the partition function.
@@ V Discussion and Conclusion
I would like to thank Mohsen Alishahiha for especial inspiration, valuable discussions, and fruitful comments. I also thank Mark Srednicki, Manuel Vielma, Julian Sonner, and Mahdieh Piranaghl for fruitful discussions. I especially acknowledge Junyu Lin for his precious help with the code at an early stage.

### v2

@@ Abstract
Recent investigations of Rényi entanglement entropy in the SYK chain of Majorana fermions have indicated that the model exhibits slow thermalization when initialized in certain states. The extent to which the heavy modes–believed to underlie this behavior–affect other aspects of the model remains an open question. In this work, I study thermalization and scrambling of information in individual energy eigenstates of the SYK chain using exact diagonalization. I show that two-point correlation functions in finite-energy eigenstates closely match their thermal counterparts and that information scrambling occurs efficiently within these pure states. These results suggest that the slow thermalization observed in entanglement dynamics does not extend to all probes of thermalization and scrambling, even in pure states.
@@ I Introduction
Chaotic quantum many-body systems are generally expected to thermalize rapidly, even in the absence of external reservoirs. The eigenstate thermalization hypothesis (ETH) provides a framework for understanding this phenomenon [1, 2, 3, 4, 5]. ETH posits that for generic observables, the expectation values in individual energy eigenstates are consistent with thermal averages, implying that a pure state far from equilibrium will evolve such that its observables approach and remain close to thermal values for most of the time.
@@
The SYK chain is a one-dimensional lattice model with quenched disorder, in which each site hosts $N$ Majorana fermions subject to random four-fermion interactions [7]. Neighboring sites are coupled via random bilinear interactions. This model inherits several important features from the original, zero-dimensional SYK model [9, 10, 11], including local criticality, an extensive zero-temperature entropy, and maximal chaos [12, 10, 13, 14]. It also displays additional features such as diffusive energy transport and a butterfly velocity that characterizes the spatial spread of chaos [7]. From a condensed matter perspective, the SYK chain is a rare example of a solvable, strongly interacting, chaotic lattice model–offering a tractable platform for studying thermalization, entanglement propagation, many-body localization, and quantum transport.
From the viewpoint of holography, the SYK chain has been conjectured to be dual to an incoherent black hole [7]. Since the SYK chain is a closed, unitary system, understanding its thermalization dynamics may yield insights into black hole formation and evaporation, and thereby inform broader discussions surrounding the black hole information paradox [15, 16].  [^1] For recent developments in resolving the information paradox, see [17, 18, 19, 20, 21, 22, 23, 24, 25].
@@ II Background
The SYK chain model [7] is a higher-dimensional generalization of the original Sachdev-Ye-Kitaev (SYK) model [28, 9, 10]. In this section, I first provide a brief overview of the original SYK model before describing the structure and properties of the SYK chain model.
@@
To introduce spatial locality, Gu et al. [7] proposed a $(1+1)$-dimensional generalization, known as the SYK chain model. This model consists of a one-dimensional lattice of $M$ SYK sites (also referred to as ``links"), with each site containing $N$ Majorana fermions and interacting with its nearest neighbors via quartic interactions. The structure is illustrated in Fig. 1.
@@ III Computational Setup
In my earlier work [8], I examined the eigenstate thermalization hypothesis (ETH) for the SYK chain model with $M=2$, 33, and $4$. Here, I focus on the case $M=4$ to investigate additional indicators of thermalization and quantum information scrambling.
@@ IV Correlation Functions and Signatures of Quantum Chaos
While many of the correlation functions considered here have been previously analyzed in the thermal ensemble [30, 31, 32, 12, 10, 14, 33], my emphasis is on their behavior in pure states, particularly in individual energy eigenstates. From the perspective of the holographic correspondence, this translates into studying how accurately correlation functions computed in an incoherent black-hole background are approximated by correlations in pure states.
The structure and methodology of this section are inspired by Ref. [26], which investigated similar questions in the 0D complex SYK model. Here, I extend those ideas to the spatially extended SYK chain to explore how thermalization and scrambling of information manifest in the presence of locality.
@@ IV.1 The Spectral Form Factor
Figure 2 displays ${\cal S}(\beta,t)$ as a function of time for the SYK chain, averaged over 1000 disorder realizations and evaluated at inverse temperature $\beta=1.5$. The data reveals the familiar three-stage structure also observed in RMT: an initial decay (the ``dip''), followed by a linear ramp, and eventual saturation at a plateau [30, 34]. While this qualitative structure resembles RMT behavior, the quantitative features–such as the dip time and plateau time–depend on model-specific details [30].
@@ IV.2 Two-Point Correlation Function in Energy Eigenstates
Figure 3 compares $G^{n}_{c}(t)$ to the corresponding thermal correlator $G^{\beta}(t)$ across different coupling regimes. To align the energies of the eigenstate and thermal ensemble, I fix the inverse temperature at $\beta=1.5$ and select the eigenstate $|n\rangle$ whose energy $E_{n}$ most closely matches the thermal average energy $E(\beta)$. The results show excellent agreement at both early and late times. Small deviations appear at intermediate times, particularly near the end of the power-law decay regime, where $G^{n}_{c}(t)$ exhibits residual oscillations around its long-time plateau, whereas $G^{\beta}(t)$ approaches the plateau more smoothly. These discrepancies diminish with increasing Hilbert space dimension and are expected to vanish in the thermodynamic limit $N\to\infty$, which is of particular relevance to the conjectured gravitational dual, as it corresponds to the semiclassical regime.
Figure 3: Two-point correlation functions of the hopping operator between two sites of the SYK chain with $NM=24$ Majorana fermions. Each correlation function is averaged over 10001000 realizations of the disorder. The eigenstate correlators $G^{n}_{c}(t)$ and the canonical correlators $G^{\beta}(t)$ exhibit excellent agreement across all three choices of intra- and inter-site coupling strengths.
@@ IV.3 Four-Point Correlation Function in Energy Eigenstates
To align the energy of the eigenstate with that of the thermal ensemble, I fix the inverse temperature to $\beta=1.5$ and select the eigenstate $|n\rangle$ whose energy $E_{n}$ is closest to the thermal average energy,
$E(\beta)=\frac{1}{Z}\text{Tr}[e^{-\beta H}H],$ (33)
where $Z=\text{Tr}[e^{-\beta H}]$ denotes the partition function.
@@ V Discussion and Conclusion
I would like to thank Mohsen Alishahiha and Akbar Jafari for their special inspiration and insightful discussions. I am also grateful to Mark Srednicki, Manuel Vielma, Julian Sonner, and Mahdieh Piranaghl for many fruitful conversations. I especially acknowledge Junyu Lin for his invaluable assistance with the code during the early stages of this work.